# Exploring the Stability Gap in Continual Learning: The Role of the Classification Head

### v1

@@ 1 Introduction
  • We are the first work to investigate the phenomenon of stability gap in the context of different parts of the network.
@@ 2.1 Continual Learning
Continual learning aims to enable learning from changing data streams and mitigate catastrophic forgetting [27] - destructive overwriting of previously acquired knowledge that appears when learning from new data. Continual learning usually assumes that the data come in the form of sequential tasks, and once the learner finishes processing the task it can no longer access its data for training; however, it is still evaluated on the data from all the tasks seen so far. The main scenarios analyzed in the continual learning community are task-incremental learning, class-incremental learning, and domain-incremental learning [40, 26]. Task-incremental learning and class-incremental learning usually analyze a scenario where new tasks contain new classes, with task-incremental learning assuming access to task identity information during prediction and class-incremental learning assuming no access to such information, which makes it a more challenging scenario. Domain-incremental learning instead operates on the set of the same classes but with the data distribution changing between tasks (e.g. changing camera, lightning, or image style). Continual learning approaches are usually categorized as regularization-based, replay-based or architecture-based, with many works also proposing hybrid approaches [26]. Regularization-based methods try to prevent forgetting by penalizing changes to model weights [16] or activations [22]. Replay- or exemplar-based methods enforce the stability of the network by including a small subset of old data during training on a new task to rehearse previously learned knowledge; the exemplars can be either stored samples from old tasks [32, 6], samples generated by a generative model [36] or stored features from the old data [3]. Architectural methods, or parameter isolation methods, try to alleviate forgetting by allocating specific network parameters to specific tasks or growing the network while training [33, 21, 41, 44]. Most of the more modern methods try to combine the strengths of multiple approaches and propose hybrid solutions [30, 14, 46, 1]
@@ 3.2 Metrics
where $f_{t}^{n}$ refers to a model obtained after $n$ gradient updates on a task $t$. This metric gives us an insight into how much the model degrades during training. However, note that it does not take into account final accuracy, and a learner whose performance initially drops to zero and then recovers all its accuracy on an old task throughout the training on a new task will still exhibit a high stability gap. To properly measure and compare the stability gap in cases where the initial model already exhibits very low accuracy, normalize the metric with the starting accuracy.
@@ 4.6 Task Recency Bias
In Fig. 8 we present task confusion matrices after training on the final task. We observe severe task recency bias for finetuning – predictions of the classes from previous tasks are much less frequent than predictions from the last task. For NMC, however, the bias is much less evident and the task confusion matrix is closer to the ideal case of a diagonal matrix.
To continually evaluate the network’s bias towards predicting the most recent task, we propose Latest Task Prediction Bias (LTB) metric and define it as the mean of the values from matrix $C$ for tasks $11\leq i\leq t-$ 1 mispredicted as $t$:
@@ Appendix A Task-Aware Accuracy
Appendix B Other approaches
We expand our investigation to regularization-based (LwF [22], SS-IL [1]), and parameter-isolation (EWC [16]) methods. We use a constant memory buffer for all the methods (2000 exemplars).
B.1 Task-Agnostic results
@@
B.2 Latest-Task Prediction Bias
@@
Figure 14: More advanced methods yield improved latent representations, while NMC further alleviates issues related to the classifier head. Observations from main experiments scale to other approaches.
@@
Appendix C Different CNN architectures

### v2

@@ 1 Introduction
  • We are the first to investigate the stability gap in the context of different parts of the network.
@@ 2.1 Continual Learning
Continual learning aims to enable learning from changing data streams and mitigate catastrophic forgetting [27] - destructive overwriting of previously acquired knowledge that appears when learning from new data. Continual learning usually assumes that the data come in the form of sequential tasks, and once the learner finishes processing the task it can no longer access its data for training; however, it is still evaluated on the data from all the tasks seen so far. The main scenarios analyzed in the continual learning community are task-incremental learning, class-incremental learning, and domain-incremental learning [40, 26]. Task-incremental learning and class-incremental learning usually analyze a scenario where new tasks contain new classes, with task-incremental learning assuming access to task identity information during prediction and class-incremental learning assuming no access to such information, which makes it a more challenging scenario. Domain-incremental learning instead operates on the set of the same classes but with the data distribution changing between tasks (e.g. changing camera, lighting, or image style). Continual learning approaches are usually categorized as regularization-based, replay-based or architecture-based, with many works also proposing hybrid approaches [26]. Regularization-based methods try to prevent forgetting by penalizing changes to model weights [16] or activations [22]. Replay- or exemplar-based methods enforce the stability of the network by including a small subset of old data during training on a new task to rehearse previously learned knowledge; the exemplars can be either stored samples from old tasks [32, 6], samples generated by a generative model [36] or stored features from the old data [3]. Architectural methods, or parameter isolation methods, try to alleviate forgetting by allocating specific network parameters to specific tasks or growing the network while training [33, 21, 41, 44]. Most of the more modern methods try to combine the strengths of multiple approaches and propose hybrid solutions [30, 14, 46, 1]
@@ 3.2 Metrics
where $f_{t}^{n}$ refers to a model obtained after $n$ epochs on a task $t$. This metric gives us an insight into how much the model degrades during training. However, note that it does not take into account final accuracy, and a learner whose performance initially drops to zero and then recovers all its accuracy on an old task throughout the training on a new task will still exhibit a high stability gap. To properly measure and compare the stability gap in cases where the initial model already exhibits very low accuracy, we normalize the metric with the starting accuracy.
@@ 4.6 Task Recency Bias
In Fig. 8 we present task confusion matrices after training on the final task. We observe severe task recency bias for finetuning with a linear head – predictions of the classes from previous tasks are much less frequent than predictions from the last task. For NMC, however, the bias is much less evident and the task confusion matrix is closer to the ideal case of a diagonal matrix.
To continually evaluate the network’s bias towards predicting the most recent task, we propose the Latest Task Prediction Bias (LTB) metric and define it as the mean of the values from matrix $C$ for tasks $11\leq i\leq t-$ 1 mispredicted as $t$:
@@ 5 Conclusions
Acknowledgments
We gratefully acknowledge Poland’s high-performance Infrastructure PL-Grid for providing computer facilities and support within computational grant no. PLG/2024/017385. This research was partially funded by National Science Centre, Poland, grant no: 2022/45/B/ST6/02817.
@@ Appendix A Task-Aware Accuracy
Appendix B Continual evaluation metrics
The stability gap (SG) is a per-task metric, making it particularly useful for analyzing the dynamics of a model’s behavior across individual tasks during sequential learning. Similarly, worst-case accuracy (WC-ACC) and average minimum accuracy (min-ACC) are metrics that can be evaluated on a per-iteration basis, providing further insights into the model’s performance during training. Visualizing these metrics for all tasks on the plots is more informative than considering only the final scores, as it highlights how the model evolves and adapts over time. We present the plots of those metrics in Fig. 14.
@@ Appendix B Continual evaluation metrics
Appendix C Other approaches
We expand our investigation to regularization-based (LwF [22], SS-IL [1]), and parameter-isolation (EWC [16]) methods. We use a constant memory buffer for all the methods (2000 exemplars). We also present continual evaluation metrics on the whole training on CIFAR100 (see Fig. 18Fig. 19, and Fig. 20).
C.1 Task-Agnostic results
@@
C.2 Latest-Task Prediction Bias
@@
Figure 14: Fine-tuning on CIFAR100 (10 tasks). Continual evaluation metrics through full training of 10 tasks.
Figure 15: TAg accuracy. More advanced methods yield improved latent representations, while NMC further alleviates issues related to the classifier head. Observations from main experiments scale to other approaches.
Figure 16: For completeness we also present TAw accuracy scores.
@@
Figure 18: LwF on CIFAR100 (10 tasks). Continual evaluation metrics.
Figure 19: EWC on CIFAR100 (10 tasks). Continual evaluation metrics.
Figure 20: SS-IL on CIFAR100 (10 tasks). Continual evaluation metrics.
Appendix D Different CNN architectures